Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 124     Entry time: 11 Jun 2024, 00:43
Author: Abhisakh Sarma 
Group: SPB 
Subject: Shift summary 

**X-ray delivery**

  • 7.0 keV, up to 202 pulses at 564 kHz, 1.0 mJ.

**Optical laser delivery**

  • ns laser for jet illumination

**Achievements/Observations**

  • Data collection completed for all the requested samples from user.


**Issues**

  • Run repeater behave once abnormally.  Run repeater started with 30 minutes duration, but the first run (run #238) stopped after 40 seconds and next run (run #239) started.

 

ELOG V3.1.4-7c3fd00